# UniRVQA: A Unified Framework for Retrieval-Augmented Vision Question Answering via Self-Reflective Joint Training

### v1

@@ 4.3. Ablation Study
Additionally, we analyze the effectiveness of main components in our proposed framework to answer the following research questions. All experiments are conducted on the OK-VQA dataset.
@@
RQ1: How does joint training on a unified framework work? To investigate how does joint training on a unified framework affects performance, we construct two variants of UniRVQA by separating the training process of retriever and answer generator in different ways. We keep the late interaction and reflective-answering mechanism in all variants. The results are summarized in Table 5. Specifically, $\mathrm{UniRVQA^{-}}$ still uses a unified framework which will first be trained on the retrieval task, followed by training on the answer generation task. $\mathrm{UniRVQA^{--}}$ employs the more traditional setting of two separate models as the retriever and the generator. By training the model separately as a retriever and a generator, $\mathrm{UniRVQA^{--}}$ performs relatively well in all metrics, although slightly worse than the full model. This indicates that base MLLMs have the potential to handle both knowledge-intensive tasks separately. In the proposed unified framework, the retriever and answer generator further share the same network, reducing model size significantly and allowing both tasks to complement each other, resulting in improved performance.
When inspecting the performance from $\mathrm{UniRVQA^{-}}$, the results show that simply unifying the framework without proper training strategy design leads to a significant drop in performance. This suggests that naive multistage training may hinder the model’s ability to share parametric knowledge effectively. The better performance on UniRVQA verifies our assumption that by sharing parametric knowledge across tasks, the model can better leverage capabilities learned from one task to improve the performance on the other.
@@
RQ2: How can the Reflective-Answering enhance the system performance? To investigate the impact of reflective-answering mechanism, we conduct ablation studies where the mechanism is disabled during the training and inference stage respectively. The results, presented in Fig. 4, show that removing reflective-answering leads to a significant drop in performance compared to the complete UniRVQA model. Specifically, removing self-reflection during training results in the most substantial decrease, with a reduction of over 6.91% in VQA score and more than 7.5% in EM for both base models. This highlights the critical role of self-reflection in improving model effectiveness during training. In addition, turning off self-reflection during inference hinders performance by approximately 1.3%, further demonstrating that reflective-answering is crucial not only for learning but also for making more accurate predictions. Therefore, encouraging the model to wisely rely on its implicit knowledge during inference yields accuracy gains ranging from 1.22% to 3.03%. We attribute the effectiveness of our proposed reflective-answering mechanism to its ability to alleviate the model’s dependence on external knowledge. This mechanism reminds the model to leverage its own implicit knowledge, which has been empirically proven to be useful (Yang et al., 2022; Gui et al., 2021). With reflective-answering, the model is more discerning in its use of external information, avoiding reliance on irrelevant data when it already possesses sufficient implicit knowledge to answer accurately.
RQ3: How does the external knowledge support the proposed system? Here, we further investigate the contributions of external knowledge in our experiments firstly by removing them during the inference, as in Figure 4. We conclude that removing the external knowledge support will directly lead to on average 4.1% reductions in answer accuracy, which confirm the necessity of external knowledge when answering knowledge-intensive questions. Compared the standard fine-tuning method that provide the model with final answer as training signals (Table 2, line 15 and 17), UniRVQA framework brings about 2% improvement in answering accuracy. Specifically, UniRVQA (InstructBLIP) improves the fine-tuned counterparts from 60.47% to 62.64% in EM and from 55.50% to 57.38% in VQAScore. Such improvement indicate that incorporating the external knowledge during the training process appropriately can also be beneficial to enhance the reasoning ability of answer generator. The theoretical mechanism is also worth to be investigated in the future research.
@@ 4.4. Case Studies
The right column features two examples from InfoSeek. The top-right example asks for the specific location where the facility is standing, a question that would be very hard to answer without particular external knowledge. This also highlights the complexity of questions in InfoSeek. UniRVQA accurately retrieves an encyclopedic document describing the ferris wheel installed on the “Place de la Concorde” in Paris, showcasing the model’s ability to effectively identify and use fine-grained information to answer challenging questions. We also note that although the base model was not able to provide the precise location, it could still identify the city in the image, which indicates that the base MLLM contains some fundamental knowledge that can be potentially leveraged. The bottom-right failure case is challenging. The model struggled with identifying the species of the black bird in the image, however still managed to retrieve generally relevant information about “blackbird”. We emphasize that a potential area for improvement is the fine-grained entity retrieval, particularly in distinguishing visually similar entities.
5. Conclusion and future work
In this paper, we propose a Unified Retrieval-Augmented Vision Question Answer framework (UniRVQA), which can effectively adapts general-purpose MLLMs for both fine-grained retrieval and knowledge-intensive answer generation tasks, through a self-reflective joint training framework and incorporating a reflective-answering mechanism to optimize implicit and explicit knowledge utilization. Extensive experiments demonstrate that UniRVQA can elevate underperforming base models to leading positions. UniRVQA also achieve competitive answering and retrieval performance, compared to state-of-the-art models, all while maintaining a smaller model size and being training efficient—making KB-VQA research more accessible. The ablation study shows that the unified framework can enhance the performance by sharing parametric capabilities to complement tasks along RAG.
In the future, we aim to enhance system performance by focusing on more accurate entity retrieval and recognition. One key direction will involve refining the model’s ability to identify and extract relevant entities from large knowledge bases, as this plays a critical role in generating contextually precise answers. The other direction lies in identifying the mechanism that external knowledge works to help with the life cycle of Knowledge-intensive VQA systems.

### v2

@@ 4.3. Ablation Study
We analyze the effectiveness of the main components in our proposed framework to answer the following research questions. The ablation study is conducted on the OK-VQA dataset.
@@
RQ1: How well does collaborative parametric knowledge calibration perform? To answer the question, we construct two variants of our model by separating the training process of retriever and answer generator in different ways. We keep the late interaction and reflective-answering mechanism in all variants. The results are summarized in Table 4. Specifically, $\mathrm{UniRVQA^{--}}$ employs the traditional setting of separating the retriever and the generator. By training the models separately, $\mathrm{UniRVQA^{--}}$ performs worse than our proposed $\mathrm{UniRVQA}$. Compared to the setting where two MLLMs handle knowledge-intensive tasks separately, our proposed unified framework enables retriever and answer generator to share the same network, reducing the model size significantly and allowing both tasks to complement each other, resulting in improved performance.
Additionally, $\mathrm{UniRVQA^{-}}$ uses a unified framework but without collaborative knowledge calibration, where the network is first trained on the retrieval task, followed by the answer generation task. The results show that simply unifying the framework without proper training strategy design leads to a significant drop in performance. This suggests that naive multistage training may hinder the model’s ability to share parametric knowledge effectively. The better performance on UniRVQA verifies our assumption that by performing collaborative parametric knowledge calibration, the model can better leverage capabilities learned from one task to improve the performance on the other.
@@
RQ2: How can the Reflective-Answering enhance the system performance? To investigate the impact of reflective-answering mechanism, we conduct experiments by disabling the mechanism during the training and inference stage respectively. The results presented in Fig. 4, show that removing reflective-answering leads to a significant drop in performance compared to the complete UniRVQA model. Specifically, removing self-reflection during training results in the most substantial decrease, with a average reduction of over 6.91% in VQA score and more than 7.5% in EM for both base models. This highlights the critical role of self-reflection in improving model performance during training. Additionally, turning off self-reflection during inference hinders performance by 1.3% on average, demonstrating that reflective-answering is crucial not only for learning but also for making more accurate predictions. Specifically, incorporating the reflective-answering mechanism into inference yields accuracy gains ranging from 1.22% to 3.03%. We attribute the effectiveness of our proposed mechanism to its ability to alleviate the model’s dependence on external knowledge. With reflective-answering, the model is more discerning in its use of external information, avoiding reliance on irrelevant data when it already possesses implicit knowledge to answer accurately.
RQ3: How does the fine-grained knowledge learning (late-interaction) support the system? Here, we further investigate the contributions of fine-grained knowledge in our experiments by removing them during the inference. As shown in Figure 4, removing the external knowledge support will directly lead to an average reduction of 4.1% in answer accuracy, which confirms the necessity of fine-grained external knowledge processing when answering knowledge-intensive questions. Compared to the standard fine-tuning method where only the final answer is provided as training labels (Table 1, line 15 and 17), the UniRVQA framework improves answer accuracy by 2% on average. Specifically, with the retrieval augmentation turned off, UniRVQA (InstructBLIP) improves the fine-tuned counterparts from 60.47% to 62.64% in EM and from 55.50% to 57.38% in VQAScore. Such improvement indicates that incorporating the fine-grained knowledge learning during the training process appropriately can be beneficial to the reasoning ability of the answer generator.
@@ 4.4. Case Studies
The right column features two examples from InfoSeek. The top-right example asks for the specific location where the facility is, a question that is hard to answer without particular external knowledge. This also highlights the complexity of questions in InfoSeek. UniRVQA accurately retrieves an encyclopedic document describing the ferris wheel installed on the “Place de la Concorde” in Paris, showcasing the model’s ability to effectively identify and use fine-grained information to answer challenging questions. We also note that although the base model was not able to provide the precise location, it could still identify the city in the image, which indicates that the base MLLM contains some fundamental knowledge that can be potentially leveraged. The bottom-right failure case is challenging. The model struggled with identifying the species of the black bird in the image, however still managed to retrieve generally relevant information about “blackbird”. We emphasize that a potential area for improvement is the fine-grained entity retrieval, particularly in distinguishing visually similar entities.
5. Conclusion
In this work, we present a unified retrieval-augmented KB-VQA framework with collaborative parametric knowledge calibration to address the limitations of decoupled retrieval and generation stages. By enabling parametric knowledge sharing between retriever and generator, and integrating both a late-interaction mechanism for fine-grained multimodal understanding and a reflective-answering mechanism for knowledge boundary assessment, our training framework significantly improves the adaptability of general MLLMs to knowledge-intensive tasks. Extensive experiments demonstrate that our approach not only achieves state-of-the-art performance but also substantially boosts the capabilities of base MLLMs, which offers a practical and efficient solution for future KB-VQA research.